### v1

# InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions

Abstract
Understanding and designing biomolecules, such as proteins and small molecules, is central to advancing drug discovery, synthetic biology, and enzyme engineering. Recent breakthroughs in Artificial Intelligence (AI) have revolutionized biomolecular research, achieving remarkable accuracy in biomolecular prediction and design. However, a critical gap remains between AI’s computational power and researchers’ intuition, using natural language to align molecular complexity with human intentions. Large Language Models (LLMs) have shown potential to interpret human intentions, yet their application to biomolecular research remains nascent due to challenges including specialized knowledge requirements, multimodal data integration, and semantic alignment between natural language and biomolecules. To address these limitations, we present InstructBioMol, a novel LLM designed to bridge natural language and biomolecules through a comprehensive any-to-any alignment of natural language, molecules, and proteins. This model can integrate multimodal biomolecules as input, and enable researchers to articulate design goals in natural language, providing biomolecular outputs that meet precise biological needs. Experimental results demonstrate InstructBioMol can understand and design biomolecules following human instructions. Notably, it can generate drug molecules with a 10% improvement in binding affinity and design enzymes that achieve an ESP Score of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This highlights its potential to transform real-world biomolecular research.
Introduction
Understanding and designing biomolecules is fundamental to natural science research. Biomolecules, such as proteins and small molecules, play essential roles in biological processes, and their precise manipulation is key to advancements in drug discovery, synthetic biology, and enzyme engineering [1, 2, 3]. Recent Artificial Intelligence (AI) breakthroughs have transformed research in these areas [4, 5]. Tools like AlphaFold3 [4] and RoseTTAFold All-Atom [5] have revolutionized biomolecular structure prediction, offering unprecedented accuracy and speed. Despite these advancements, a crucial challenge persists: how to effectively understand biomolecules using natural language and design them according to human intentions. This presents a gap between AI’s computational capacity and researchers’ needs to apply them to real-world problems. Consider the scenario of a researcher tasked with designing a new drug to target a protein involved in a complex disease. Traditionally, this process involves navigating vast amounts of biomolecular data, interpreting biological and chemical relationships, and iterating through trial-and-error to engineer molecules with specific properties. While AI has enhanced many aspects of this workflow, current tools often struggle to align molecular complexity with intuitive, human-driven goals articulated in natural language.
To fully unlock the potential of AI in biomolecular science, there is an urgent need for systems that can seamlessly bridge the gap between biomolecular data and human intention. Such a system would allow researchers to articulate their design goals in natural language and receive molecular outputs that meet precise biological needs—whether it be generating a drug molecule tailored to a specific protein or designing an enzyme optimized for a chemical process. This alignment of AI with human expertise and intuition is crucial for advancing research in areas where creativity and expert knowledge play a central role.
Notably, Large Language Models (LLMs) [6] have demonstrated their impressive capacity to understand human intention and generate human-like responses helpfully and safely [7, 8]. This capability derives from its vast number of learnable parameters, training on expansive corpora, and strong alignment with human preferences. Nevertheless, despite their potential to comprehend and design biomolecules in natural language following human intentions, leveraging LLMs for biomolecular research is still in its nascent stage and presents several challenges. Firstly, biomolecular research demands substantial specialized knowledge. General-purpose LLMs may not possess deep insights into this field because they are not tailored for specialized domains. Consequently, the lexical and semantic gap between the natural language and the language used to describe biomolecules presents its challenges. For instance, “C” could represent an alphabet letter in English, a carbon atom in a chemical molecule, or cysteine in the context of amino acids in proteins. Such semantic discrepancies might confuse models that are not specifically designed for these contexts. Secondly, unlike the primarily textual focus of general LLMs, biomolecules are intrinsically multimodal. For example, molecules are often represented by sequences like SMILES [9] or SELFIES [10]. Also, molecules can be inherently depicted as 2D graphs, featuring atoms as nodes and chemical bonds as edges, or as 3D structures, noting the spatial arrangements of atoms. Similarly, proteins are described through FASTA [11] sequences for their amino acid composition, and they also have 3D structures, essential for understanding their interactions and functions in space [12]. The effective utilization of this multimodal data presents a unique challenge for text-focused LLMs. Lastly, and most importantly, general LLMs struggle to align human intention in biomolecular tasks. To achieve the desired performance in a specific domain, LLMs need to be trained with alignment to acquire task-specific knowledge and patterns [13]. Instruction plays a critical role in this alignment process. By providing carefully curated, domain-specific instructions, LLMs are guided to develop a deeper understanding and more precise execution of specific tasks. In particular, to follow instructions for biomolecules, mastering the alignment between natural language, molecules, and proteins is essential. Complex tasks such as designing molecules for target proteins or creating enzymes for substrates necessitate simultaneous processing of natural language, molecules, and proteins.
Although several recent endeavors [14, 15, 16, 17, 18, 19] have sought to tailor the LLMs for biomolecular tasks via extensive instructions, they encounter two primary limitations: (1) They typically align natural language with either molecules or proteins, but not both, lacking any-to-any alignment. (2) They also fall short in processing multimodal biomolecules, failing to align the multimodal data with natural language.
Figure 1: An overview of InstructBioMol. a, InstructBioMol is a unified multimodal Large Language Model for natural language, molecules, and proteins. It can accept inputs in the form of natural language text, multimodal molecule and protein data and generate outputs as natural language, molecules, or proteins in the textual form. b, The Motif-Guided Multimodal Feature Extraction Module processes 2D graphs and 3D structures of molecules, as well as 1D sequences and 3D structures of proteins. Pre-trained modality-specific encoders obtain representations from these inputs, which are then processed by a Transformer Encoder. The Transformer Decoder, using motif prompts and learnable queries, produces multimodal features for integration into the language model. c, We collect datasets on a hundred-million scale, categorized into continual pretraining data and instruction-tuning data. Instruction-tuning is used to achieve an any-to-any alignment between molecule, protein, and natural language. A two-stage instruction-tuning paradigm enables the model to learn from low-quality extensive data (stage-1) to high-quality refined data (stage-2). d, InstructBioMol achieves alignment between molecules and natural language, proteins and natural language, as well as molecules and proteins. This enables it to follow human instructions, facilitating the understanding and design of biomolecules.
In this study, we introduce InstructBioMol, which exhibits the following key characteristics:
  • Biomolecular instruction following. InstructBioMol integrates natural language and biomolecules within one Large Language Model, becoming the first to achieve any-to-any pairwise alignment between natural language, molecules, and proteins. By leveraging a curated hundred-million scale instruction-tuning dataset, the model is empowered to understand and design biomolecules according to human intention.
  • Multimodal data understanding. We propose a motif-guided multimodal feature extraction module. It utilizes pre-trained encoders to capture various features, including 2D topological and 3D geometric details of molecules, and 1D sequence and 3D geometric properties of proteins. Also, we design a motif prompt extractor, which leverages biological knowledge embedded in motifs to guide the multimodal feature fusion.
  • Serving as a research copilot and supporting practical biomolecular tasks . The value of InstructBioMol is its role as a digital research assistant, supporting researchers in biomolecular studies and discoveries. It excels in employing natural language processing to explore biomolecules, such as answering questions related to protein functions or designing novel molecules based on textual descriptions. Moreover, InstructBioMol demonstrates its potential in solving practical tasks, such as drug discovery and enzyme design.
In the experiments, InstructBioMol is thoroughly assessed for its proficiency in understanding and designing molecules and proteins following human instructions, showing its ability to align natural language with biomolecules. Specifically, InstructBioMol achieves an overall improvement of 12% in understanding and designing molecule and protein benchmark tasks. Additionally, we explore the model’s application in generating drug molecules aimed at specific target proteins and designing enzyme proteins for particular substrates. Experimental results reveal that drug molecules designed by InstructBioMol exhibit an improvement of 10% in binding affinity, while the enzymes it designs achieve an ESP Score [20] of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. These exercises confirm InstructBioMol’s applicability in practical biomolecular research scenarios. Overall, these results not only validate the broad generalization of InstructBioMol as a Large Language Model that bridges natural language and biomolecules, but also underscore its potential for a wide range of applications within the life sciences.
Results
Overview of InstructBioMol
The overview of InstructBioMol is presented in Figure 1a. InstructBioMol is a unified multimodal Large Language Model that simultaneously handles natural language, molecules, and proteins. It accepts inputs in natural language or multimodal molecules and proteins, and generates natural language, molecules, or proteins in textual form. To process multimodal data of both molecules and proteins, we develop a Motif-Guided Multimodal Feature Extraction Module (Figure 1b). This module employs a Transformer Encoder-Decoder structure [21]. It encodes 2D graphs and 3D structures for molecules and 1D sequences along with 3D structures for proteins. Using pre-trained lightweight encoders, the multimodal inputs are encoded into corresponding representations, and subsequently fed into a Transformer Encoder. In the Transformer Decoder, we extract the biological knowledge in the motifs using a motif prompt extractor, which serves as guiding information for the fusion of multimodal features. The fused features are then integrated into the language model. For model training, outlined in Figure 1c, a wide range of data is collected, comprising a continual pretraining dataset and an instruction-tuning dataset. The continual pretraining dataset comprises molecules, proteins and natural language texts derived from scientific literature. The instruction-tuning dataset consists of data pairs between natural language, molecules, and proteins. The training process is divided into two stages. First, continual pretraining is employed to augment domain-specific knowledge in the field of biomolecular scientific research. Then, instruction-tuning is performed to achieve an any-to-any pairwise alignment among natural language, molecules, and proteins. We employ a staged instruction-tuning pipeline, learning from large-scale data (stage-1) to refined data (stage-2) to gradually improve performance. As a result, InstructBioMol aligns natural language, molecules, and proteins in an any-to-any manner, demonstrating competency across a broad spectrum of biomolecular tasks, as shown in Figure 1d. This includes solving practical challenges, like the discovery of molecule drugs for target proteins, and the design of enzymes for specific substrates, following human intention.
InstructBioMol can understand and design molecules following human intention
Experimental Setup.
We evaluate InstructBioMol’s capability in understanding and designing molecules through two tasks: (1) molecule captioning, which involves generating a textual description for a molecule; and (2) description-based molecule generation, where a molecule is generated based on a provided textual description. These tasks are introduced by ref. 14, and use the ChEBI dataset [22], which contains molecules and their corresponding descriptions including structure, function, origin, etc. Following ref. 14, for the first task, we select BLEU [23], ROUGE [24], and METEOR [25] as evaluation metrics. For the second task, we employ EXACT to measure the exact match of generated molecules, and using a range of similarity metrics — including BLEU, LEVENSHTEIN [26], MACCS FTS [27], RDK FTS [28], MORGAN FTS [29], and FCD [30] — to assess the similarity between generated molecules and ground truth. Furthermore, we utilize VALIDITY to evaluate the chemical validity of the generated molecules. Details of evaluation metrics are described in Methods. In this experiment, we evaluate two types of baselines. (1) Generalist language models. We assess two variants of general-purpose models. First, GPT-3.5 (zero-shot), a commonly used Large Language Model. Second, GPT-3.5 (10-shot MolReGPT) and GPT-4 (10-shot MolReGPT), which are adaptations of GPT-3.5 and GPT-4, respectively, using MolReGPT’s few-shot in-context learning approach [31]. (2) Molecule-specific enhanced language models. These models are further fine-tuned for molecular tasks, building on generalist models. Baselines include ChemDFM [32], InstructMol [33], MolT5 [14], BioT5 [16], and BioT5+ [18]. Notably, InstructMol is incapable of molecule generation.
Results.
Quantitative results on molecule captioning and molecule generation are in Table 1 and Table 2, respectively. The performance of molecule-specific models significantly surpasses that of generalist language models, primarily due to the finetuning of the latter on domain-specific instruction datasets. This finetuning facilitates an effective alignment between natural language and chemical molecular knowledge. These findings suggest that general language models lack sufficient expertise in specialized domains, which can be effectively compensated by leveraging instruction-tuning. Notably, the experimental results demonstrate that InstructBioMol performs best across almost all evaluation metrics. Specifically, for the molecule captioning task, InstructBioMol yields an average improvement of 0.9% across all metrics. In description-based molecule generation task, the exact match accuracy (EXACT) of generated molecules increases by 0.7%. Furthermore, an average improvement of 2.0% is observed in molecular fingerprint similarity metrics (MACCS FTS, RDK FTS, and MORGAN FTS). These results indicate that InstructBioMol exhibits higher accuracy and efficacy in both understanding and generating chemical molecular information. We attribute this success to the extensive use of high-quality instruction data, which enables the model to comprehensively align molecules and natural language and achieve superior performance across molecular tasks. Some examples in Supplementary Information Section 4.1 provide a detailed analysis of the results generated by InstructBioMol.
Table 1: Performance comparison on molecule captioning task. (↑) / (↓) denotes a higher / lower value is better. The best performance is marked as bold.
BLEU-2 (↑) BLEU-4 (↑) ROUGE-1 (↑) ROUGE-2 (↑) ROUGE-L (↑) METEOR (↑)
GPT-3.5 (zero-shot) 10.3 5.0 26.1 8.8 20.4 16.1
GPT-3.5 (10-shot MolReGPT) 56.5 48.2 62.3 45.0 54.3 58.5
GPT-4 (10-shot MolReGPT) 60.7 52.5 63.4 47.6 56.2 61.0
ChemDFM 32.1 26.5 49.0 37.4 48.3 40.2
InstructMol 47.5 37.1 56.6 39.4 50.2 50.9
MolT5 64.4 57.2 70.8 58.4 65.3 68.1
BioT5 63.5 55.6 69.2 55.9 63.3 65.6
BioT5+ 66.6 59.1 71.0 58.4 65.0 68.1
InstructBioMol 66.3 59.3 72.0 60.1 66.8 69.1
Table 2: Performance comparison on description-based molecule generation task. (↑) / (↓) denotes a higher / lower value is better. The best performance is marked as bold.
BLEU (↑) EXACT (↑) LEVENSHTEIN (↓) MACCS FTS (↑) RDK FTS (↑) MORGAN FTS (↑) FCD (↓) VALIDITY (↑)
GPT-3.5 (zero-shot) 48.9 1.9 52.13 70.5 46.2 36.7 2.05 80.2
GPT-3.5 (10-shot MolReGPT) 79.0 13.9 24.91 84.7 70.8 62.4 0.57 88.7
GPT-4 (10-shot MolReGPT) 85.7 28.0 17.14 90.3 80.5 73.9 0.41 89.9
ChemDFM 83.9 43.2 16.90 90.1 82.9 75.9 - 97.6
MolT5 85.4 31.1 16.07 83.4 74.6 68.4 1.20 90.5
BioT5 86.7 41.3 15.10 88.6 80.1 73.4 0.43 100.0
BioT5+ 87.2 52.2 12.78 90.7 83.5 77.9 0.35 100.0
InstructBioMol 87.7 52.9 13.65 91.8 85.8 80.5 0.24 99.0
Figure 2: Model performance on protein understanding and design benchmarks. a, Regarding protein understanding, models are evaluated by examining their performance in answering questions related to protein family, subcellular location, name, and function. The evaluation metrics employed include BLEU-2, BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, and METEOR. b, For description-based protein generation task, the accuracy and biological validity of the generated proteins are assessed using metrics including Identity, Alignment, BLOSUM Substitution, and Validity.
InstructBioMol can understand and design proteins following human intention
Experimental Setup.
We evaluate the model’s ability to understand and design proteins in the following two tasks: (1) answering questions about the properties of proteins, including protein family, subcellular location, official name, and function; (2) generating protein sequences based on the textual descriptions. For the protein property answering task, we apply the same evaluation metrics as for the molecule captioning task to assess the similarity between the generated answers and ground truth. For the description-based protein generation task, we use Identity, Alignment, and BLOSUM Substitution to measure the similarity of the generated proteins to ground truth. In particular, Identity captures exact amino acid matches; Alignment assesses the similarity of subsequences between proteins; and BLOSUM Substitution evaluates potential evolutionary relevance through amino acid substitutions. Higher scores on these three metrics indicate that the generated sequence is more similar to the ground truth. We also report the Validity for evaluating the biological validity of protein generation. These metrics are described in Methods in detail. In the experiments, we compare the following baseline methods: (1) the general-purpose language model GPT-3.5. GPT-3.5 is evaluated in three variants: GPT-3.5 (zero-shot), GPT-3.5 (5-shot, random), and GPT-3.5 (5-shot, similarity). The GPT-3.5 (5-shot, random) and GPT-3.5 (5-shot, similarity) variants utilize an in-context learning paradigm (detailed in Supplementary Information), where the former randomly selects 5 examples from the training set as prompts, and the latter selects the 5 most similar examples to the given query. (2) Protein-specific enhanced language models, which extend general language models to protein-related tasks, including Mol-Instructions [17], InstructProtein [15], BioT5+ [18], ProtT3 [34], BioMedGPT [19], and ProteinDT [35]. However, among these models, BioMedGPT and ProtT3 cannot generate protein sequences, while ProteinDT cannot answer protein property-related questions using natural language.
Results.
Quantitive results on answering protein properties and description-based protein generation are in Figure 2a and Figure 2b, respectively. Based on the experimental results, we have reached the following conclusions: Firstly, InstructBioMol demonstrates the best performance in both tasks. In tasks related to answering questions about protein properties, InstructBioMol outperforms previous state-of-the-art (SOTA) methods by 13.1% on average. For protein generation tasks, InstructBioMol achieves 0.9%, 5.7%, and 0.510 improvements in identity, alignment, and BLOSUM substitution, respectively, compared to previous SOTAs, and comparable validity of the generated proteins. Secondly, using domain-specific instruction alignment is effective. In tasks related to answering protein property questions, InstructBioMol shows an average improvement of 80.3% compared to GPT-3.5 (zero-shot). In protein generation tasks, InstructBioMol achieved a 17.5% increase in identity and a 28.7% increase in alignment compared to GPT-3.5 (zero-shot). These results clearly demonstrate that models aligned with domain-specific instructions exhibit significantly enhanced capabilities in handling tasks within specific domains compared to general models. This underscores the importance of customized models in specialized domains, particularly in highly specialized fields like protein engineering, where the integration of domain-specific knowledge and instructions can greatly enhance the model’s practicality and accuracy. In Supplementary Information Section 4.2, we provide a detailed analysis of InstructBioMol’s outstanding performance on protein function answering tasks through some examples. Additionally, example cases on description-based protein generation task demonstrate that InstructBioMol can design de-novo proteins with high structural similarity, closely aligning with the structures of ground truth.
Figure 3: Model performance on drug discovery and enzyme design. a-c, Performance comparison on drug discovery. (a), The generated drug-like molecules are evaluated from multiple perspectives, including binding affinity (e.g., Vina Score and High Affinity), general properties (SA, QED, and Validity), and an overall evaluation metric (Success Rate). b-c, A detailed analysis of Vina Scores is presented, including (b) top-1, top-5, top-10, and all Vina Scores, as well as (c) the distribution of Vina Scores for all generated molecules. d, Performance comparison on enzyme design. Evaluation metrics include similarity metrics Identity and Alignment; interaction metrics ESP Score and Vina Score; and Validity.
InstructBioMol enables target protein-based drug discovery following human intention
Experimental Setup.
Designing molecules that can bind to specific proteins is one of the most challenging tasks in drug discovery [36]. The chemical space is vast, yet the subset of molecules with desirable biological activity is relatively small. Drug discovery typically involves searching this expansive space for molecules that can bind to specific targets, such as disease-related proteins. As a research copilot, InstructBioMol can design molecule drugs for target proteins following human intention, thereby reducing the time and cost of drug development. For experimental validation, we have established a series of evaluation metrics. Following refs.37, 38, evaluation metrics can be divided into three perspectives, including (1) quantitative estimates of binding affinity: Vina Score [39] and High Affinity. Vina Score reflects the best possible binding affinity, where a lower score indicating stronger affinity. High Affinity measures the ratio of generated molecules that bind stronger than the reference molecules in the test set. (2) General molecular properties: synthetics accessibility (SA) [40], drug-likeness (QED) [41], and Validity. (3) Overall assessment following ref. 38: Success Rate, which measures the ratio of molecules with binding affinity versus favorable properties (Vina Score $<$ -8.18, QED $>$ 0.25, SA $>$ 0.59). Details of metrics are described in Methods. To make a comparison, we take three variants of GPT-3.5 (zero-shot, 5-shot random, and 5-shot similarity) and DrugGPT [42] as baselines, and also take the molecules in the test set as the Reference baseline. 100 target proteins are selected in the test set, and for each target protein, 100 molecules are generated.
Results.
Figure 3a presents the experimental results. where InstructBioMol demonstrates superior performance across three key dimensions: binding affinity, general properties, and overall assessment. Specifically, it improves High Affinity and Success Rate by 25.9% and 21.9%, respectively, compared to previous state-of-the-art (SOTA) methods. This highlights InstructBioMol’s enhanced capability in generating drug-like molecules with high affinity for target proteins and favorable intrinsic properties. Additionally, it achieves an outstanding generation validity of 99.9%. Figure 3b illustrates the average Vina Scores for top-1, top-5, top-10, and all generated molecules. It can be observed that InstructBioMol consistently outperforms other methods under these settings, and is the only approach where the average scores for all generated molecules surpass the reference values. This suggests that the quality of the molecules generated by InstructBioMol is comparable to the ground truth in the dataset. Moreover, InstructBioMol proves to be the most effective method for designing molecules with the best Vina Scores for most target proteins. (As shown in Figure 4a in Supplementary Information, InstructBioMol achieves the best performance on 35% of the targets.) Figure 3c shows the distribution of Vina Scores for all generated molecules. The distribution reveals that molecules generated by InstructBioMol have a lower mean and reduced variance, further confirming that the overall quality of these molecules is superior to that of other methods.
Figure 4: Results of ablation analysis on the impact of training strategy and the multimodal data integration. a, The impact of removing specific training stages is evaluated across four tasks: molecule captioning, description-based molecule generation, protein function answering, and description-based protein generation. The models compared include InstructBioMol and its two variants: one variant retains the continual-pretraining stage but removes the stage-1 instruction-tuning (w/ continual-pretraining, w/o stage-1 instruction-tuning), while the other variant removes both the continual-pretraining stage and the stage-1 instruction-tuning (w/o continual-pretraining, w/o stage-1 instruction-tuning). b, The comparison of multimodal data integration, involving InstructBioMol and its various variants: w/o 1D encoder, w/o 2D encoder, and w/o 3D encoder, which correspond to the removal of specific modality inputs. Additionally, w/o motif denotes the removal of motif prompts, and w/o multimodal represents the removal of the entire Motif-Guided Multimodal Feature Extraction Module.
InstructBioMol enables target substrate-based enzyme design following human intention
Experimental Setup.
Enzymes, as biological catalysts, can accelerate chemical reactions in various biological processes [43]. In enzymatic reactions, substrates are the small molecules that are catalytically converted by enzymes. By binding to specific substrates and acting upon them, enzymes significantly enhance the conversion efficiency of the substrates. Designing enzymes that can bind to specific substrates is a crucial and challenging research problem. InstructBioMol can assist researchers in designing protein enzymes for specific substrates, thereby advancing the progress of efficient enzyme design. To establish experimental validation, we split a test set containing 100 substrates, and for each target substrate, 100 enzyme proteins are generated. The following evaluation metrics are employed, which are divided into three groups: (1) similarity between generated proteins and ground truth, including Identity and Alignment. (2) Substrate-enzyme interaction, assessed using ESP Score [20] and Vina Score. ESP Score measures the interaction capability of the designed enzymes to their substrates, with a range of 0-100, where higher scores indicate stronger interactions. Vina Score quantifies the strength of binding affinity between the designed enzymes and substrates, with lower scores indicating higher affinity. (3) Validity is used to assess whether generated enzymes are biologically valid. Details of metrics are in Methods. For comparison, we employ three variants of GPT-3.5 (zero-shot, 5-shot random, 5-shot similarity) and RFdiffusionAA [5] as baseline models.
Results.
The performance of the generated protein enzymes across various evaluation metrics is presented in Figure 3d. InstructBioMol demonstrates the best performance in terms of similarity to ground truth, interaction capability with substrates, and exhibits superior generation validity. Specifically, InstructBioMol achieves improvements of 13.3 in ESP Score and 0.7 in Vina Score, indicating a stronger potential for substrate binding compared to baseline methods. Notably, InstructBioMol attains an ESP Score of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This demonstrates that enzymes designed by InstructBioMol can bind their corresponding substrates with high affinity. Figure 4b in Supplementary Information further analyzes the top-1 ESP Score of the proteins generated for each substrate, revealing that InstructBioMol achieves the best performance on 66% of the substrates. Additionally, Supplementary Figure 4c presents the top-1 Vina Score on each substrate, with InstructBioMol attaining the best performance on 89% of the substrates. These findings suggest that InstructBioMol holds significant potential in generating highly efficient and specific protein enzymes, offering more effective solutions for fields such as biocatalysis.
Ablation analysis
Analysis of Training Strategy.
We first analyze the impact of different training strategies on model performance. Figure 4a presents a comparison of InstructBioMol with two of its variants: one variant retains continual pretraining and stage-2 instruction-tuning but removes stage-1 instruction-tuning (denoted as "w/ continual-pretraining, w/o stage-1 instruction-tuning"), while the other variant only retains stage-2 instruction-tuning but removes both continual pretraining and stage-1 instruction-tuning (denoted as "w/o continual-pretraining, w/o stage-1 instruction-tuning"). From these comparisons, we derive the following conclusions:
Firstly, in tasks related to molecule understanding and design (such as molecule captioning and description-based molecule generation), it is observed that the contribution of continual pretraining is relatively minor, whereas the significance of stage-1 instruction-tuning is more pronounced. In the molecule captioning task, removing stage-1 instruction-tuning results in an average performance drop of 7.3%, while further removal of continual pretraining leads to an additional drop of only 1.2%. In the molecule generation task, removing stage-1 instruction-tuning decreases the exact accuracy of generated molecules by 12.4%, and further removal of continual pretraining caused a subsequent drop of only 1.1%. We hypothesize that this is due to the critical role played by the molecular IUPAC name data used in stage-1 instruction-tuning, which effectively bridges and aligns molecular structure with natural language. Secondly, in description-based protein generation task, we find that contribution of continual pretraining is more significant, while the impact of stage-1 instruction-tuning is comparatively smaller. Specifically, removing stage-1 instruction-tuning leads to a decrease in the BLOSUM Substitution score by 0.187, and further removal of continual pretraining results in an additional drop of 0.326. This indicates that proteins, with their complex and intrinsic relationships such as sequence homology, benefit greatly from the evolutionary information captured by the broad protein sequence dataset used in the continual pretraining stage, thereby significantly enhancing the quality of the generated proteins.
Analysis of Multimodal Data Integration.
Next, we explore the impact of incorporating multimodal data on model performance. As shown in Figure 4b, we compare InstructBioMol and several of its variants on molecule captioning and protein function answering tasks. These variants include the removal of the 2D encoder (w/o 2D encoder) and 3D encoder (w/o 3D encoder) for molecular data, as well as the removal of the 1D encoder (w/o 1D encoder) and 3D encoder (w/o 3D encoder) for protein data. Additionally, we consider the removal of motif prompts (w/o motif) and the entire Motif-Guided Multimodal Feature Extraction Module (w/o multimodal). For comparison, the InstructBioMol variant used here excludes both the continual-pretraining stage and stage-1 instruction-tuning, serving as a basis for the ablation study on multimodal inputs.
The analysis results indicate that removing the Motif-Guided Multimodal Feature Extraction Module leads to a significant decline in model performance, with an average decrease of 4.7% and 7.1% on molecular and protein-related tasks, respectively. These findings underscore the importance of multimodal feature extraction in these tasks. Specifically, multimodal feature extraction can integrate information from different modalities, such as molecular structures and protein structures, thereby providing more comprehensive and accurate feature representations, which compensate for the limitations of single-modal features. Further analysis reveals that 3D data has a minimal impact on molecular-related tasks, with the removal of the 3D encoder resulting in only a 0.5% performance loss. Notwithstanding, for protein-related tasks, the influence of 3D data is more pronounced, with the removal of the 3D encoder causing a 4.2% decline in performance. This may be because the 3D information of molecules primarily describes the spatial distribution of atoms, overlooking detailed information about chemical bonds and functional groups. In contrast, the structure of proteins is fundamental to their functional performance, serving as a critical determinant of protein interactions and the formation of active sites. In conclusion, although different modalities contribute variably to different tasks, overall, multimodal feature extraction enhances the model’s performance in molecule and protein-related tasks. These results emphasize the importance of integrating multimodal data in biomolecular tasks.
Discussion
In this study, we propose InstructBioMol, a multimodal Large Language Model capable of following human instructions for understanding and designing biomolecules. To address the limitation that general-purpose language models cannot handle multimodal biomolecular data, we design a motif-guided multimodal feature extraction module. This module extracts multimodal features from biomolecules and leverages the knowledge embedded in motifs to guide the fusion of these features, which are then integrated into the language model. During training, we employ a training paradigm that involves “continual pretraining followed by instruction-tuning”, based on extensive pretraining and instruction-tuning data. In instruction-tuning, we adopt a staged strategy to progressively reduce the data size while enhancing data quality. Through comprehensive instruction-tuning, InstructBioMol becomes the first model capable of achieving any-to-any alignment between natural language, molecules, and proteins. Our experiments demonstrate the effectiveness of these alignments across a range of tasks involving natural language, molecules, and proteins. InstructBioMol is not only capable of understanding and designing molecules or proteins following human intention, but it can also design drug-like molecules for target proteins or enzyme catalysts for reaction substrates. This indicates InstructBioMol’s potential as a research copilot, offering valuable insights and inspiration to researchers, with practical applications in drug and enzyme design.
One limitation of InstructBioMol lies in the constraints imposed by computational resources, preventing it from fully supporting all biomolecules, such as DNA and RNA. Furthermore, it has not been comprehensively trained across all biomolecular tasks, which limits its ability to handle certain additional tasks, such as chemical reaction prediction. However, based on the current model architecture and training framework, InstructBioMol exhibits strong extensibility. By incorporating more multimodal encoders and expanding its vocabulary, it can enhance its encoding and generation capabilities for other biomolecules, and it can be easily adapted to new tasks through additional instruction data. Another concern is the profound implications and potential risks associated with the integration of Large Language Models and biomolecules. Ensuring alignment between LLMs and human ethics is crucial. For instance, when utilizing LLMs to design novel biomolecules, adherence to strict ethical guidelines is essential to avoid irresponsible experimentation and potential biosafety hazards. In the future, we plan to enhance the alignment of InstructBioMol with human values and ethics, ensuring its consistency with societal norms and enabling it to inspire biomolecular innovations safely and effectively.
We believe the core value of InstructBioMol lies in pioneering a new paradigm for processing biomolecular data using Large Language Models, showcasing the potential of general intelligence in handling diverse tasks in one model. With the increase in computational resources, the enrichment of training data, and the enhanced alignment with human ethics, InstructBioMol is expected to evolve and support a broader range of tasks effectively and safely, laying the groundwork for advancing Artificial General Intelligence (AGI) in scientific research.
Methods
Model Architecture
The architecture of InstructBioMol (Figure 1a) consists of two components: the Motif-Guided Multimodal Feature Extraction Module (Figure 1b) and the Biomolecular Vocabulary-expanded Language Model. The former is designed to extract multimodal features of biomolecules, while the latter handles a unified processing of textual natural language, molecule and protein data, as well as the extracted multimodal features. Specifically, in the Motif-Guided Multimodal Feature Extraction Module, we employ lightweight frozen pre-trained encoders to extract features from each modality separately, and leverage the biological knowledge embedded in motifs to guide the fusion of these multimodal features. Within the Biomolecular Vocabulary-expanded Language Model, to mitigate potential interference among data from different domains, we expand the vocabulary to accommodate molecules and proteins, and standardize the input format for their multimodal features.
Motif-Guided Multimodal Feature Extraction Module
Biomolecules exhibit inherent multimodality, characterized by diverse sequential and structural representations across various domains [44, 45, 46, 12]. This complexity cannot be fully captured by any single modality in isolation. In this module (Figure 1b), we incorporate 2D-graph and 3D-structure for molecules, alongside 1D-sequence and 3D-structure for proteins to leverage the multitude of perspectives available. 2D-graph of molecules highlights the basic skeleton, while 3D-structure provides insights into molecular docking and interaction. For proteins, 1D-sequence delineates the fundamental arrangement of amino acids, and 3D-structure unlocks understanding of functional sites and foldings. We utilize frozen pre-trained encoders to process each modality separately, and then leverage the inherent biological knowledge within motifs to guide multimodal feature fusion, which enhances comprehension and processing of complex biological data.
Multimodal Inputs and Encoders.
For molecules, the 2D-graph modality is defined as $m_{2D}=(V,E)$, where $V$ stands for atomic nodes and $E$ represents chemical bonds between these atoms. The 3D-structure modality is defined as $m_{3D}=(V,\mathbf{C})$, with $V$ indicating a set of atoms and $\mathbf{C}\in\mathbb{R}^{|V|\times 3}$ representing the spatial coordinates of these atoms. We leverage a pre-trained 5-layer GIN [47, 48] as the 2D-graph encoder $f_{m}^{2D}$, and Geoformer [49] as the 3D-structure encoder $f_{m}^{3D}$, to derive the respective modality inputs’ representations:
$\mathbf{H}^{2D}_{m}=f_{m}^{2D}(m_{2D}),\quad\mathbf{H}^{3D}_{m}=f_{m}^{3D}(m_{3D}),$ (1)
where $\mathbf{H}^{2D}_{m}\in\mathbb{R}^{|V|\times d_{m}^{2D}}$ and $\mathbf{H}^{3D}_{m}\in\mathbb{R}^{|V|\times d_{m}^{3D}}$ are the obtained molecular 2D and 3D representations, respectively. For proteins, the 1D-sequence modality is characterized by $p_{1D}=(s_{1},s_{2},...,s_{N})$, where each $s_{i}$ is an amino acid. And the 3D-structure modality can be represented as $p_{3D}=(S,\mathbf{C})$, where $S$ is the amino acid sequence and $\mathbf{C}\in\mathbb{R}^{N\times 4\times 3}$ denotes the coordinates of four backbone atoms (N, C, CA, O) in each amino acid. Here, we adopt pre-trained ESM2-35M [46] and SaProt-35M [50] as the 1D-sequence encoder $f_{p}^{1D}$ and 3D-structure encoder $f_{p}^{3D}$, respectively, to encode the two modalities of proteins:
$\mathbf{H}^{1D}_{p}=f_{p}^{1D}(p_{1D}),\quad\mathbf{H}^{3D}_{p}=f_{p}^{3D}(p_{3D}),$ (2)
where $\mathbf{H}^{1D}_{p}\in\mathbb{R}^{N\times d_{p}^{1D}}$ and $\mathbf{H}^{3D}_{p}\in\mathbb{R}^{N\times d_{p}^{3D}}$ represent the obtained protein 1D and 3D representations, respectively.
Motif Prompt Extractor.
In molecules, a motif often represents a functional group or substructure, playing a crucial role in determining molecular function and structure [51, 52]. Similarly, in proteins, motifs are sequences of consecutive amino acids carrying specific biological functions, forming foundational elements for protein functionality [53]. To integrate the essential prior knowledge within motifs, we introduce a motif prompt. The motif prompt is designed to highlight key regions within biomolecules that are pivotal for understanding their function and interaction. By acting as a conditional input, it guides the multimodal feature extraction process towards features that are relevant to the identified motifs, thereby increasing the biological relevance of the extracted features. Specifically, we denote the motifs in a molecule as $\mathbf{T}_{m}=[t_{1}^{m},t_{2}^{m},...,t_{N_{m}}^{m}]$, and motifs in a protein as $\mathbf{T}_{p}=[t_{1}^{p},t_{2}^{p},...,t_{N_{p}}^{p}]$, where $t_{i}\in\{0,1\}$ indicates the presence ($t_{i}$=1) or absence ($t_{i}$=0) of the $i$-th motif in the molecule or protein, and $N_{m}$ and $N_{p}$ represent the total counts of predefined motifs in molecules and proteins, respectively. Subsequently, the motif prompt for molecules or proteins is computed as:
$\mathbf{P}_{m}=\mathbf{T}_{m}\mathbf{M}_{m},\quad\mathbf{P}_{p}=\mathbf{T}_{p}\mathbf{M}_{p},$ (3)
where $\mathbf{P}_{m}\in\mathbb{R}^{d}$ and $\mathbf{P}_{p}\in\mathbb{R}^{d}$ are motif prompt of molecule and protein, respectively. $\mathbf{M}_{m}\in\mathbb{R}^{N_{m}\times d}$ and $\mathbf{M}_{p}\in\mathbb{R}^{N_{p}\times d}$ are two learnable matrices. In detail, we obtain $\mathbf{T}_{m}$ by computing the Functional-Class FingerPrint (FCFP) [29] of a molecule with a radius of 2 and a length of 1024. The motifs of protein are collected from the UniProt [54] database, and details are described in Supplementary Information.
Joint Multimodal Feature Extraction.
The intrinsic complexity of biomolecules necessitates a modeling approach that is not only capable of capturing the detailed nuances of each modality but also adept at discerning the intricate interrelationships between them. Based on the extracted representations of a single modality, we propose to exploit the complementarities and redundancies between the various modalities to discover potentially more informative and robust representations than those obtained from any single-modality data. Herein, a Transformer [21] Encoder-Decoder architecture is employed for effective extraction and fusion of features from these heterogeneous sources. The reason for the choice of the Transformer architecture is that the inherent self-attention mechanism allows for the dynamic weighting of various parts of the input data, enabling to focus on the most relevant features across and within modalities. And the flexibility of the Encoder-Decoder facilitates the cross-modal integration of features, with the encoder capturing the salient features and the decoder synergizing those to construct a fused multimodal representation.
Firstly, the single-model representations obtained from Eq. 1 and Eq. 2 are transformed and then concatenated:
$\begin{split}&{\mathbf{H}^{2D}_{m}}^{\prime}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{H}^{2D}_{m})),\quad{\mathbf{H}^{3D}_{m}}^{\prime}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{H}^{3D}_{m})),\quad\mathbf{H}_{m}=[{\mathbf{H}^{2D}_{m}}^{\prime}\oplus{\mathbf{H}^{3D}_{m}}^{\prime}],\\ &{\mathbf{H}^{1D}_{p}}^{\prime}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{H}^{1D}_{p})),\quad{\mathbf{H}^{3D}_{p}}^{\prime}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{H}^{3D}_{p})),\quad\mathbf{H}_{p}=[{\mathbf{H}^{1D}_{p}}^{\prime}\oplus{\mathbf{H}^{3D}_{p}}^{\prime}],\end{split}$ (4)
where $\mathbf{H}_{m}\in\mathbb{R}^{2|V|\times d}$ and $\mathbf{H}_{p}\in\mathbb{R}^{2N\times d}$ are used as the inputs of the Transformer Encoder for molecule and protein, respectively, and $\oplus$ denotes the concatenation operation. We utilize the motif prompt obtained from Eq. 3 as the initial input to the Transformer Decoder. By directing the focus of the Transformer-Decoder toward these motifs, this approach endeavors to anchor the multimodal feature extraction in biologically significant referents. This ensures that the resultant fused features are not only data-derived but also deeply rooted in the biological realities of molecular and protein functionalities. Additionally, the input to the Transformer Decoder includes a sequence of learnable queries. Formally, the joint multimodal feature extraction is defined as:
$\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{(1+N_{q})}]=\mathrm{Transformer}(\mathrm{Enc}(\mathbf{H}),\mathrm{Dec}([\mathbf{P}\oplus\mathbf{Q}])).$ (5)
To simplify, subscripts are omitted since molecules and proteins undergo the same processing. Here, $\mathbf{P}$ and $\mathbf{H}$ are derived from Eq. 3 and Eq. 4 respectively. $\mathbf{Q}$ denotes a sequence of learnable queries as $\mathbf{Q}=[\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{N_{q}}]\in\mathbb{R}^{N_{q}\times d}$, and $\mathbf{Z}\in\mathbb{R}^{(1+N_{q})\times d}$ is the extracted multimodal features.
Biomolecular Vocabulary-expanded Language Model
Language Model Backbone.
InstructBioMol is designed to be compatible with any GPT-style [55] language model. In this study, we specifically adopt Llama-2-7B [7] for further training.
Expanding Vocabulary.
In this work, we use SELFIES [10] to represent molecules and FASTA [11] (sequence of amino acids) to represent proteins. Despite their utility, a notable conflict arises among natural language, molecules, and proteins, where identical tokens may imply entirely different meanings. For example, the token "C" in English simply refers to the letter C, but in molecular contexts, it represents a carbon atom, and in protein sequences, it denotes cysteine. This ambiguity prevents natural language vocabularies from distinguishing these entities effectively. Hence, we introduce extended vocabulary for molecules and proteins, integrating them with the original natural language vocabulary. Specifically, for molecules, we utilize the pair of brackets within SELFIES along with the meaningful group of atoms they encapsulate as a token. For instance, "[C]" denotes a Carbon atom. For proteins, we introduce a specific prefix "<p>" for each amino acid, such as "<p>C" for cysteine. Furthermore, we introduce specialized tokens to differentiate between modalities. These include "<SELFIES>", "</SELFIES>" for molecule SELFIES sequence, "<FASTA>", "</FASTA>" for protein amino acid sequence, as well as "<MOL>", "</MOL>" "<PROT>" "</PROT>" signify outputs from the Motif-Guided Multimodal Feature Extraction Module for molecules and proteins, respectively. This deliberate separation of different modalities ensures the preservation of each modality’s intrinsic integrity and prevents model confusion regarding the meanings of different modalities.
Input Formation.
By expanding the vocabulary and incorporating diverse multimodal features, we integrate molecules or proteins into textual formats, thereby augmenting the language model’s capacity to interpret biomolecules. Specifically, we concatenate these multimodal features $[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{(N_{q}+1)}]$ obtained from Eq. 5 with sequence-modality input, and label them with special tokens:
[…]
$x_{multimodal\_mol}$ and $x_{multimodal\_prot}$ denote molecules and proteins, respectively. The inputs also encompass instructions $x_{insuction}$ in natural language, e.g., "What is the function of this protein", as well as description text $x_{text}$, e.g., "The molecule is a member of benzenes, a sulfone and a member of triazoles.". The composition of the inputs adapts based on the specific task. For instance, in generating molecular descriptions, the inputs include $x_{instruction}$ and $x_{multimodal}$, whereas in generating molecules from descriptions, the inputs consist of $x_{instruction}$ and $x_{text}$. The form of input corresponding to each task is detailed in Table 4.
Data type Entries Tokens Source
molecule 100M 4B PubChem
protein 49M 15B Uniref50
natural language 6M 8B PubMed, bioRxiv, ChemRxiv
Table 3: Statistics of continual pretraining dataset.
Data Collection
We collect datasets on a hundred-million-scale, including a continual pretraining dataset (Table 3), and an instruction-tuning dataset (Table 4). The continual pretraining dataset comprises molecules and proteins in textual format, and natural language texts derived from scientific literature, enabling the model to develop a foundational adapting to biomolecular research. The instruction-tuning dataset contains various alignment pairs: molecule-natural language, protein-natural language, and molecule-protein (Figure 1d), achieving any-to-any alignment among molecules, proteins and natural language. Additionally, both molecular and protein data in the instruction-tuning dataset are multimodal, incorporating 2-D and 3-D structures of molecules and 3-D structures of proteins.
Continual Pretraining Dataset
For molecules, we collect 115 million entries from PubChem [56], filtering out those with atomic numbers exceeding 50, resulting in 100 million entries and 4 billion tokens after tokenization. For proteins, we use the Uniref50 dataset [57], which comprises 49 million entries and 15 billion tokens after tokenization. Recognizing the limitations of general language models trained on generalized corpora lacking biomolecular insights, we augment our data collection with literature specific to the research domain. Abstracts of scientific papers are collected across several sources, including PubMed [58], bioRxiv [59], and ChemRxiv [60], to enrich our training corpus. The incorporation of literature from these repositories enhances the domain-specific knowledge of InstructBioMol. This subset includes 6 million abstracts, resulting in 8 billion tokens.
Instruction-tuning Dataset
Molecule-Natural Language Pairs.
The dataset is sourced from two databases. The first is from PubChem, where we collect molecules with the IUPAC (International Union of Pure and Applied Chemistry) [61] name. This naming convention establishes a standardized nomenclature, fostering uniformity and clarity across the chemical community. 30 million molecules with IUPAC names are sampled from the filtered set in the previous step. The second source is ChEBI [22] data from ref. 14, comprising molecules alongside their descriptions. These descriptions encapsulate various facets of molecular structure, function, synthesis methodologies, etc.
Protein-Natural Language Pairs.
Data originate from two databases: SwissProt and TrEMBL [54]. These databases provide textual descriptions of proteins, covering four key aspects: name, family, location, and function. We utilize SwissProt data collected in ref. 19 and curate TrEMBL data from the UniProt [62] database. To ensure data quality and diversity, we filter TrEMBL dataset using UniRef50, resulting in approximately 25 million proteins. Additionally, for proteins described in at least three of four aspects, we consolidate these descriptions into a comprehensive summary using ChatGPT, creating data aligned from natural language to proteins.
Molecule-Protein Pairs.
For molecule and protein pairs, we focus on two key applications: the discovery of drug-like molecules for specific target proteins and the design of enzyme proteins to catalyze specific substrates. Specifically, for generating molecules to specific target proteins, we use data from BindingDB [63] collected in ref. 64. BindingDB is a public database primarily focusing on the interactions between proteins, identified as potential targets, and small, drug-like molecular ligands. On the other hand, for designing an enzyme to catalyze a particular substrate, we draw upon data from the Rhea [65] database collected in ref. 20. Rhea is an expert-curated database of chemical reactions of biological interest, where enzyme-catalyzed reactions are curated from peer-reviewed literature.
Multimodal Data.
The extraction of multimodal features necessitates access to diverse data types relating to molecules and proteins. To accomplish this, we employ RDKit [66] to convert molecules to 2D-graph and optimize them using ETKDG [67] and Merck Molecular Force Field [68] to obtain 3D-structure. For 3D-structure of proteins, we download the predicted 3D structures from the AlphaFold Protein Structure Database [69] via the UniProt ID of each protein. This guarantees that the molecule and protein data have rich multimodal characteristics, laying a solid foundation for their application in downstream processes.
Training Strategy
We start with a pretrained language model and continue pretraining it on the continual pretraining dataset in a self-supervised causal language modeling objective [55]. Subsequently, we employ instruction-tuning, to establish an any-to-any alignment among natural language, molecules, and proteins. This involves aligning specific instructions and inputs with appropriate responses, represented as $(x_{instruction},x_{input})\rightarrow y$, where $x_{input}$ may include multimodal molecules $x_{multimodal\_mol}$, proteins $x_{multimodal\_prot}$ defined in Eq. 6, or natural language in textual format $x_{text}$, and $y$ denotes corresponding responses such as natural language $y_{text}$, molecular sequences $y_{mol}$, or protein sequences $y_{protein}$. To achieve a thorough alignment, we introduce a bidirectional alignment task for each pairwise alignment among natural language, molecules and proteins. For example, for molecule-natural language pairs, one task generates textual descriptions from molecular data: $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$, and another task generates molecules from descriptions: $(x_{instruction},x_{text})\rightarrow y_{mol}$. The instruction-tuning is optimized under a causal language modeling objective:
$\min_{\theta}\mathcal{L}_{CE}\left(\mathrm{LM}\left(x_{instruction},x_{input}\right),y\right),$ (7)
where $\mathcal{L}_{CE}$ is cross-entropy loss, $\theta$ is all the model parameters except four frozen pre-trained multimodal encoders, $\mathrm{LM}(\cdot)$ denotes the language model’s prediction, and $y$ is the label.
Two-Stage Instruction-tuning.
Despite the collection of a broad range of data, the quality of this data exhibits considerable variability across sources. For example, the ChEBI database offers a broader spectrum of molecular descriptions compared to the natural-language-like structure descriptions provided by IUPAC names in PubChem. Similarly, while data within the SwissProt undergo meticulous manual curation, entries in the TrEMBL do not benefit from such rigorous calibration. The tradeoff between the scale and quality of data poses significant challenges to model performance and generalizability. To address this issue, we adopt a two-stage instruction-tuning strategy designed to exploit the extensive data initially, then progressively direct the focus towards the insights offered by higher-quality datasets. Initially, in stage-1, the model is trained across all the available instructions. This stage leverages the diversity and volume of data to build a foundation on biomolecular alignment. Subsequently, in stage-2, the model undergoes further fine-tuning on a subset of higher-quality data. This approach harnesses both the expansive coverage of lower-quality data and the precision inherent in high-quality data, facilitating an efficient and effective utilization of the dataset. The scale and specific details of the different datasets used in the two stages are presented in Figure 1c and Table 4.
Sub-dataset Task type Scale Instruction stage-1 stage-2
Molecule - Natural Language Alignment
PubChem $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 30M Give the IUPAC name of the following molecule.
PubChem $(x_{instruction},x_{text})\rightarrow y_{mol}$ 30M Generate a molecule in SELFIES that fits the provided IUPAC name.
ChEBI $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 26K Provide a caption for the molecule below.
ChEBI $(x_{instruction},x_{text})\rightarrow y_{mol}$ 26K Generate a molecule in SELFIES that fits the provided description.
Protein - Natural Language Alignment
TrEMBL_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 25M What is the official name of this protein?
TrEMBL_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 5M What is the protein family that this protein belongs to?
TrEMBL_Locaction $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 2M What is the subcellular location of this protein?
TrEMBL_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 1M What is the function of this protein?
TrEMBL_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 2M Generate a protein matching the following description.
SwissProt_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 455K What is the official name of this protein?
SwissProt_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 370K What is the protein family that this protein belongs to?
SwissProt_Location $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 275K What is the subcellular location of this protein?
SwissProt_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 411K What is the function of this protein?
SwissProt_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 394K Generate a protein matching the following description.
Molecule – Protein Alignment
BindingDB $(x_{instruction},x_{multimodal\_prot})\rightarrow y_{mol}$ 335K Generate a drug molecule binding to the target protein.
Rhea $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{prot}$ 190K Generate an enzyme that can catalyze for the given substrate.
Table 4: Statistics of instruction-tuning dataset.
Implementation Details
We use Pytorch [70] to implement the modal. The model is trained on 8 80G NVIDIA H800 GPUs. Additionally, we adopt the DeepSpeed ZeRO-1 [71] and BF16 for computational efficiency. The total number of training steps is 1.5 million. The steps for continual pretraining, stage-1 instruction-tuning and stage-2 instruction-tuning are 600,000, 500,000, and 400,000 respectively. We use the AdamW optimizer with $(\beta_{1},\beta_{2})$ set to $(0.9,0.95)$. We follow a linear learning rate schedule, warming up from 0 to maximum learning rate 1e-5 over the first 2,000 steps, and decaying the final learning rate down to 0. During training, all parameters are trainable except for the modality encoders $f_{m}^{2D}$, $f_{m}^{3D}$, $f_{p}^{1D}$, $f_{p}^{3D}$ in Equations 1 and 2, with the total trainable parameters being 6.8B. The ratio of different datasets sampled during training is controlled using hyper-parameters, which are detailed in Supplementary Table 1.
Evaluation Tasks and Datasets
To validate the model’s capability in molecule understanding and design, two tasks are performed: molecule captioning and description-based molecule generation. The dataset we use is consistent with ref. 14. To assess the model’s ability to understand and design proteins, we conduct experiments on protein property answering and description-based protein generation tasks. The data splits used in the experiments follow ref. 19 , with 3,000 samples selected as the test set for each task. Specifically, the protein properties include family, subcellular location, name, and function. For proteins containing at least three of these properties, we use ChatGPT to combine them into a complete description, which serves as the dataset for protein design based on the description. For the task of designing molecule drugs targeting proteins, we select 100 proteins as the test set. Similarly, for the task of designing enzymes based on substrates, we select 100 enzymes as the test set. Details of datasets are provided in the Supplementary Information Section 3.
Evaluation Metrics
Evaluation Metrics for Molecule Captioning
We leverage standard natural language generation metrics such as BLEU [23], ROUGE [24], and METEOR [25] to evaluate molecule captioning following ref. 14. These metrics measure how closely the generated captions match the reference captions.
Evaluation Metrics for Description-based Molecule Generation
The following several types of metrics proposed in ref. 14 are used to evaluate the task of generating molecules from textual descriptions:
BLEU.
Similar to BLEU scores in natural language processing, the SMILES BLEU score measures the overlap between the generated molecules and the reference in SMILES strings.
EXACT.
This metric checks for exact matches between the generated and reference molecules. It provides a strict measure of accuracy.
LEVENSHTEIN.
Levenshtein distance [26] measures the number of single-character edits required to transform the generated molecules in SMILES format into the reference string. A lower Levenshtein distance indicates a closer match.
Fingerprint Metrics.
We use three types of fingerprint metrics—MACCS FTS, RDK FTS, and Morgan FTS. These use the MACCS fingerprint [27], RDK fingerprint [28], and Morgan fingerprint [29], respectively. And then calculating the Tanimoto similarity [72] between the fingerprints of the generated and reference molecules, providing a measure of how structurally similar the generated molecules are to the reference molecules.
FCD.
FCD (Fréchet ChemNet Distance) [30] compares the distributions of features derived from ChemNet between the generated and reference molecules. A lower value indicates a closer match.
Validity.
This metric assesses the percentage of generated molecules that are syntactically valid according to chemical rules.
Evaluation Metrics for Protein Property Answering
Considering that both protein property answering and molecule captioning are natural language generation tasks, we adopt the evaluation metrics of the molecule captioning task to assess this task following ref. 14, 19.
Evaluation Metrics for Description-based Protein Generation
Identity.
This metric is designed to measure the similarity between two protein sequences by calculating their percentage identity. It firstly counts the number of identical residues by comparing each corresponding residue in the reference protein $p_{ref}$ and the generated protein $p_{gen}$. Then it normalizes the number of identical residues by the sum of the lengths of both sequences. Formally,
$Identity=\frac{2\times identical\_residues}{len(p_{ref})+len(p_{gen})}\times 100.$ (8)
This formula yields a normalized value that ranges between 0 and 100, where a value of 100 indicates perfect identity, and a value of 0 indicates no identity.
Alignment.
This metric assesses the similarity between two protein sequences by leveraging the alignment scoring, which performs sequence alignment using the Smith-Waterman algorithm [73] to identify regions of alignment subsequences between the reference protein $p_{ref}$ and the generated protein $p_{gen}$. The alignment focuses on finding the highest-scoring subsequences, which allows the comparison of potentially functionally or structurally significant regions. This metric is computed based on the alignment score, and then normalized by the combined lengths of both sequences:
$Alignment=\frac{2\times alignment\_score}{len(p_{ref})+len(p_{gen})}\times 100.$ (9)
This normalization accounts for the length variations of the proteins and ensures the metric ranges between 0 and 100, with 100 indicating perfect alignment similarity and 0 indicating no alignment.
BLOSUM Substitution.
This metric calculates the similarity between the reference protein $p_{ref}$ and the generated protein $p_{gen}$ using a BLOSUM45 [74] substitution matrix-based scoring approach. This substitution matrix is commonly employed to assess the evolutionary similarity of protein sequences by providing scores for each possible pair of amino acids based on observed substitution frequencies in homologous proteins. For each pair of residues at a corresponding position, we first retrieve the substitution score from the BLOSUM45 matrix. Then, the total score, representing the cumulative similarity of all residue pairs, is normalized by the length of the two sequences:
$BLOSUM\_Substitution=\frac{2\times\sum substitution\_matrix(a,b)}{len(p_{ref})+len(p_{gen})},$ (10)
where $a$ and $b$ are amino acids from $p_{ref}$ and $p_{gen}$, respectively, and $substitution\_matrix(a,b)$ denotes the substitution score for the pair. When $substitution\_matrix(a,b)>0$, it indicates that the substitution of one amino acid for another occurs more frequently in related proteins than would be expected by chance, suggesting conservative substitutions and likely preserving protein structure and function. On the other hand, when $substitution\_matrix(a,b)<0$, it signifies that the substitution is less common, implying a disruptive effect on protein function or structure.
Validity.
This metric is employed to evaluate the valid proportion of the generated proteins, assessing whether the generated proteins are composed of amino acid sequences.
Evaluation Metrics for Target Protein-based Drug Discovery
For drug discovery, we evaluate the generated molecules from three perspectives following ref. 37, 38:
Target Binding Affinity.
Binding affinity reflects the interaction strength between the generated molecules and the target protein. Vina Score is used to estimate this affinity, with lower scores indicating stronger binding. Specifically, we first retrieve protein structures from AlphaFold Protein Structure Database [69], then use DiffDock-L [75] to estimate the protein-molecule complex structures. Qvina [39] is then employed to compute the scores. Additionally, we introduce the High Affinity metric to measure the proportion of generated molecules that achieve better binding scores than reference molecules within the test set.
Molecular Property.
General molecular properties, such as QED (Quantitative Estimation of Drug-likeness) [41] and SA (Synthetic Accessibility) [40], are utilized to evaluate the drug-likeness and synthetic accessibility of molecules. QED provides an assessment of a molecule’s potential as a drug by considering parameters like molecular weight, lipophilicity, and polar surface area, with scores ranging from 0 to 1; higher scores indicate greater drug-likeness. SA quantifies the ease of molecule synthesis, also on a scale from 0 to 1, with higher scores reflecting simpler synthetic processes. Furthermore, Validity is employed to determine the proportion of generated molecules that are syntactically valid according to chemical rules.
Overall Assessment.
Following ref. 38. we use the Success Rate to assess the quality of the generated molecules by considering multiple factors, including binding affinity, drug-likeness, and synthetic accessibility. A molecule is successful if it meets specific thresholds for Vina Score, QED, and SA (Vina Score $<$ -8.18, QED $>$ 0.25, SA $>0.59$).
Evaluation Metrics for target Substrate-based Enzyme Design
In the task of enzyme design, since generations are protein sequences, we choose to use Indentity and Alignment metrics to assess the similarity between the generated and reference proteins. To assess the interaction between enzyme proteins and substrates, we employ two metrics: Vina Score and ESP Score [20]. Vina Score is used to quantify the strength of the interaction between the designed enzyme and its substrate. A lower value indicates a stronger interaction. Specifically, when calculating the Vina Score, we use DiffDock-L to obtain the complex structure and then use Qvina to obtain the corresponding score. ESP Score is another metric for evaluating enzyme-substrate interactions. This score ranges from 0 to 100, with higher scores indicating stronger interactions. We evaluate the model’s optimal performance by calculating the average top-1 Indentity, Alignment, Vina score, and ESP score for all substrate-specific designed proteins. Additionally, we use the Validity metric to evaluate the biological validity of the generated proteins.
Supplementary Information for
InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions
1 Implementation of InstructBioMol
1.1 Collection of Protein Motifs
We download data from the SwissProt database [54] and collect all subsequences annotated as motifs. After filtering, we retain only those subsequences that appear at least twice, resulting in a total of 4712 subsequences. These subsequences are regarded as protein motifs. For ease of processing, the presence of a motif in a given protein amino acid sequence is determined by checking whether the sequence contains the corresponding subsequence.
1.2 Training Details
During continual pretraining, we set batch size to 32 and fix sequence length to 512 tokens. To ensure balanced training on molecule, protein, and natural language data, the sampling ratio for these three types of data is fixed at 1:1:1. In instruction-tuning, we use a batch size of 24, with the maximum sequence length set to 512. By setting hyperparameters, we control the sampling ratio of different types of instruction data. The sampling ratios for the datasets used in stage-1 and stage-2 of instruction-tuning are detailed in Table 1.
1.3 Inference Settings
For inference on downstream tasks, we load the model using the bfloat16 data format. The specific inference hyperparameters for each task are detailed in Table 2.
2 Baselines
To validate the effectiveness of InstructBioMol, we conduct experiments comparing it with various baseline models. In molecule captioning and description-based molecule generation tasks, the selected baseline models include MolT5 [14], BioT5 [16], BioT5+ [18], MolReGPT [31], InstructMol [33] and ChemDFM [32]. For comparisons on other tasks, baselines are categorized into two groups. The first group comprises pre-trained models, including Mol-Instructions [17], InstructProtein [15], ProtT3 [34], BioMedGPT [19], ProteinDT [35], DrugGPT [42] and RFdiffusionAA [5]. These models are all downloaded from their official repositories and evaluated on the test set. The second group consists of baseline models constructed by us using the general-purpose Large Language Model GPT-3.5. It includes three variants: zero-shot, 5-shot random, and 5-shot similarity. In the zero-shot setting, we directly pose task-specific questions to the GPT-3.5 model. In the 5-shot random setting, five examples from the training set are randomly selected as in-context demonstrations for each test entry. In the 5-shot similarity setting, the in-context learning paradigm is also adopted, but the demonstrations are required to be the five most similar examples from the training set relative to the query. The method for computing similarity depends on the data type of the input query: when query is in natural language, TF-IDF with cosine similarity is used as the text similarity measure; for protein sequence queries, MMseq2 [76] is employed to calculate protein similarity; and for molecule queries, molecular fingerprint similarity [29, 72] is used. The specific input format for the employed GPT baseline is detailed in Table 3.
3 Datasets
3.1 Examples of Datasets
We provide the example entries of continual pretraining dataset in Table 4, and the example entries of instruction-tuning dataset in Table 5, Table 6 and Table 7.
3.2 Split of Datasets
To evaluate the model’s performance, we divide the dataset used for the stage-2 instruction tuning into training and test sets. For the dataset aligning molecules with natural language, we adopt the data split from ref. 14. For the dataset aligning proteins with natural language, we use the training data defined in ref. 19 and randomly select 3000 samples from the test set as our evaluation data. Statistics of the above two datasets are in Table 8. For the dataset aligning molecules with proteins, we account for the specific nature of the tasks. In the task of generating drug-like molecules for proteins, a single protein typically corresponds to multiple molecules. Conversely, in the task of generating enzymes for substrate molecules, a single substrate often corresponds to multiple proteins. Accordingly, we split the dataset as follows: for the former task, we select 100 target proteins and their corresponding molecules as the test set. For the latter task, we select 100 target substrates and their corresponding proteins as the test set. The specific sizes of each dataset split are detailed in Table 9.
4 Additional Results
4.1 Case Analysis for Experimental Results of Understanding and Designing Molecules
For the task of description-based molecule generation, several case examples of the ground truth and generation are presented in Figure 1. Overall, InstructBioMol demonstrates a relatively accurate analysis of molecular structure, function, origin, etc. In the task of description-based molecule generation, InstructBioMol is capable of generating molecules that are completely consistent with the ground truth in certain cases, such as molecules PubChem-CID-5281294 and PubChem-CID-31284 shown in Figure 2. This demonstrates the strong molecular design capability of InstructBioMol. However, in some other cases, the molecules generated by InstructBioMol show some discrepancies with the ground truth. Our analysis suggests that this may be due to inadequate handling of certain functional groups. For example, for PubChem-CID-118429016 and PubChem-CID-123953, the model omits certain functional groups (a hydroxyl group and a phosphate group, respectively). For PubChem-CID-179394, the model generates a chemically atypical P(O)(O)(O)O group. Overall, InstructBioMol exhibits a high level of accuracy in molecule generation tasks and shows potential for applications in fields such as drug discovery, and further optimization may be required in practical applications.
4.2 Case Analysis for Experimental Results of Understanding and Designing Proteins
In the protein function answering task, InstructBioMol generates results that closely resemble the ground truth for certain cases, such as Q9NRY2 and P73070 in Table 10. Furthermore, we observe that in some cases, the generated descriptions tend to be more detailed. For example, in the case of Q9Y2G3, the generated description includes detailed information on vesicle formation, lipid signal molecule uptake, and the establishment of the thrombopoietin gradient in platelets. Similarly, for Q9FY89, the generated description provides a more comprehensive explanation of the formation of multivesicular bodies (MVBs), specifically describing the formation mechanism of intraluminal vesicles (ILVs) within MVBs, including the invagination and scission of the endosomal membrane. It further elaborates on the function of MVBs, such as transporting their contents to lysosomes for the degradation of membrane proteins, receptors, lysosomal enzymes, and lipids. For P0CP67, the generated description offers more detailed and in-depth functional information, identifying the specific targets of the protein’s action (e.g., components of AP-1, c-Jun, and ATF2) and potential biological processes involved (e.g., regulation of circadian clock). Although InstructBioMol may provide researchers with deeper insights, further experimental validation is necessary to confirm these findings.
For the task of text-based protein generation, we present two examples in Figure 3. For the ground truth proteins, we utilize the protein structures predicted by AlphaFold Protein Structure Database [69]. For the generated protein sequences, we predict their structures using ColabFold [77]. Besides sequence similarity metrics Identity, Alignment, and BLOSUM Substitution, we also compare structural similarity metrics: TM-Score [78] and LDDT [79]. The results demonstrate that InstructBioMol is capable of de-novo design of proteins, with the designed proteins exhibiting a high degree of structural similarity to ground truth. This suggests that InstructBioMol holds significant potential in designing proteins tailored to specific functional descriptions, acting as an effective copilot to assist researchers in protein design.
Table 1: Sampling ratios of different types of instruction data during instruction-tuning. Note that in practice, the sampling ratios are scaled proportionally to ensure that the sum of the ratios for all data equals 1.
Sub-dataset Task type Sampling ratio
stage-1 stage-2
Molecule - Natural Language Alignment
PubChem $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 0.1 -
PubChem $(x_{instruction},x_{text})\rightarrow y_{mol}$ 0.1 -
ChEBI $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 0.001 0.1
ChEBI $(x_{instruction},x_{text})\rightarrow y_{mol}$ 0.001 0.1
Protein - Natural Language Alignment
TrEMBL_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 -
TrEMBL_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 -
TrEMBL_Locaction $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 -
TrEMBL_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 -
TrEMBL_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 0.1 -
SwissProt_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 0.1
SwissProt_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 0.1
SwissProt_Location $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 0.1
SwissProt_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 0.05 0.1
SwissProt_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 0.1 0.2
Molecule – Protein Alignment
BindingDB $(x_{instruction},x_{multimodal\_prot})\rightarrow y_{mol}$ 0.05 0.1
Rhea $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{prot}$ 0.05 0.1
Table 2: Inference hyperparameters for downstream tasks.
Downstream Task Inference Hyperparameter
Molecule captioning $num\_beams=5$
Description-based molecule generation $num\_beams=5$
Protein property answering $top\_p=0.1,temperature=1$
Description-based protein generation $top\_p=0.9,temperature=0.8$
Target protein-based drug discovery $top\_p=1,temperature=1$
Target substrate-based enzyme design $top\_p=0.9,temperature=0.8$
Table 3: In-context learning examples of the GPT baseline across different tasks, Where XX represents data-specific natural language descriptions, protein sequences, or molecular sequences. In the few-shot setting, the input to the GPT model consists of a template, in-context demonstrations, and a question; in the zero-shot setting, the input consists of a template and a question.
Template In-context Demonstration Question
Protein property answering
You are a biologist. Given the protein sequence, your task is to generate a family of this protein using your experienced knowledge. Please strictly follow the format, no other information can be provided. Protein sequence: XX; Protein family: XX. … Protein sequence: XX; Protein family: XX Protein sequence: XX; Protein family:
You are a biologist. Given the protein sequence, your task is to generate a subcellular localization of this protein using your experienced knowledge. Please strictly follow the format, no other information can be provided. Protein sequence: XX; Protein subcellular localization: XX. … Protein sequence: XX; Protein subcellular localization: XX Protein sequence: XX; Protein subcellular localization:
You are a biologist. Given the protein sequence, your task is to generate a name of this protein using your experienced knowledge. Please strictly follow the format, no other information can be provided. Protein sequence: XX; Protein name: XX. … Protein sequence: XX; Protein name: XX Protein sequence: XX; Protein name:
You are a biologist. Given the protein sequence, your task is to generate a function of this protein using your experienced knowledge. Please strictly follow the format, no other information can be provided. Protein sequence: XX; Protein function: XX. … Protein sequence: XX; Protein function: XX Protein sequence: XX; Protein function:
Description-based protein generation
You are a biologist. Given the protein description, your task is to design a new protein matching the description using your experienced knowledge. You MUST reply using a sequence of the capitalized initial letters of 20 amino acids and DO NOT reply with others. Please strictly follow the format, no other information can be provided. Protein description: XX, Protein sequence: XX. … Protein description: XX, Protein sequence: XX. Protein description: XX, Protein sequence:
Target protein-based drug discovery
You are a biologist. Given the protein, your task is to design a drug molecule binding to this protein using your experienced knowledge. You should only reply with SMILES and DO NOT reply with others. Please strictly follow the format, no other information can be provided. Protein sequence: XX; Molecule: XX. … Protein sequence: XX; Molecule: XX. Protein sequence: XX; Molecule:
Target substrate-based enzyme design
You are a biologist. Given the molecule, your task is to design an enzyme protein that can catalyze for this substrate using your experienced knowledge. You MUST reply using a sequence of the capitalized initial letters of 20 amino acids and DO NOT reply with others. Please strictly follow the format, no other information can be provided. Molecule: XX; Protein sequence: XX. … Molecule: XX; Protein sequence: XX. Molecule: XX; Protein sequence:
Table 4: Examples of the continual pertaining data.
Data type Data
Molecule $][C$][C][S][C][C][C][C]
Protein MLSKNNNRELKRKMEEKQDRFTIKKLSVGVASVLLGSFIMGTQAVQTAHASDDNTEDATVNSAQ NTTMEQVVPLTASTS
Natural language Antiphospholipid syndrome (APS) and heparin-induced thrombocytopenia (HIT) are thrombotic disorders due to specific autoimmune-mediated antibodies. Catastrophic APS (CAPS), also known as Asherman’s syndrome, is a life-threatening severe form of APS. Diagnostic criteria for CAPS include the development of a thrombotic event of three or more organs in less than a week with the presence of antiphospholipid antibodies and …
Table 5: Examples of the instruction-tuning data on molecule-natural language alignment.
Dataset Instruction Input Output
PubChem Give the IUPAC name of the following molecule. [C][=C][Branch1][C][F][C] [Branch1][C][O][C][C][C] 2-fluorohex-1-en-3-ol
Generate a molecule in SELFIES that fits the provided IUPAC name. 2-fluorohex-1-en-3-ol [C][=C][Branch1][C][F][C] [Branch1][C][O][C][C][C]
ChEBI Provide a caption for the molecule below. [C][C][C][C][C][O][C] [=Branch1][C][=O][C][=C][C] [=C][C][=C][Ring1][=Branch1] [C][=Branch1][C][=O][O] The molecule is a phthalic acid monoester obtained by formal condensation of one of the carboxy groups of phthalic acid with the hydroxy group of pentanol. It has a role as a xenobiotic metabolite, an anti-estrogen and a rat metabolite. It derives from a pentan-1-ol.
Generate a molecule in SELFIES that fits the provided description. The molecule is a phthalic acid monoester obtained by formal condensation of one of the carboxy groups of phthalic acid with the hydroxy group of pentanol. It has a role as a xenobiotic metabolite, an anti-estrogen and a rat metabolite. It derives from a pentan-1-ol. [C][C][C][C][C][O][C] [=Branch1][C][=O][C][=C] [C][=C][C][=C][Ring1] [=Branch1][C][=Branch1][C] [=O][O]
Table 6: Examples of the instruction-tuning data on protein-natural language alignment.
Dataset Instruction Input Output
TrEMBL_Name What is the official name of this protein? MFRRGYAKYCFDNGISIYDISLSMGHSN INTTVSYINKNSDDISIYKIFNQI Tyr recombinase domain-containing protein
TrEMBL_Family What is the protein family that this protein belongs to? MRKLMALCALAGVVLVTGCNTMAGAGKD IEKGGEKVQGAAESVKQKM Belongs to the EcnA/EcnB lipoprotein family.
TrEMBL_Locaction What is the subcellular localization of this protein? LNMAENSCIDRCVSKYWQVTNLVGQLLG NNQPPM Mitochondrion inner membrane Peripheral membrane protein Intermembrane side
TrEMBL_Function What is the function of this protein? MFDQATKLHFRGARIWLAVVEDLMAKGM RHAENVRNTLNILSTCSLL Hydrolyzes acetyl esters in homogalacturonan regions of pectin. In type I primary cell wall, galacturonic acid residues of pectin can be acetylated at the O-2 and O-3 positions. Decreasing the degree of acetylation of pectin gels in vitro alters their physical properties.
TrEMBL_Description Generate a protein matching the following description. The protein is Phospholipid scramblase. It belongs toPhospholipid scramblase family. FUNCTION: It may mediate accelerated ATP-independent bidirectional transbilayer migration of phospholipids upon binding calcium ions that results in a loss of phospholipid asymmetry in the plasma membrane. MQEMLTDADTFSATFPLNLDVN VKAGLLAATFLIDFLYFEDE
SwissProt_Name What is the official name of this protein? DCCRKPFRKHCWDCTAGTPYYGYSTRNI FGCTC Mytimycin.
SwissProt_Family What is the protein family that this protein belongs to? TRSGGACNSHNQCCDDFCSTATSTCV Belongs to the conotoxin O1 superfamily.
SwissProt_Location What is the subcellular localization of this protein? MRIAKIGVIALFLFMALGGIGGVMLAGY TFILRAG Cell inner membrane; Single-pass membrane protein.
SwissProt_Function What is the function of this protein? GKIPIGAIKKAGKAIGKGLRAVNIASTA HDVYTFFKPKKRH Has antibacterial activity against Gram-positive and Gram-negative bacteria.
SwissProt_Description Generate a protein matching the following description. The protein is Photosystem II reaction center protein M, PSII-M. The protein is located in the plastid, specifically on the chloroplast thylakoid membrane. It is a single-pass membrane protein. It belongs to the PsbM family. The protein is one of the components of the core complex of photosystem II (PSII). PSII is a light-driven water:plastoquinone oxidoreductase that uses light energy to abstract electrons from H(2)O, generating O(2) and a proton gradient subsequently used for ATP formation. It consists of a core antenna complex that captures photons, and an electron transfer chain that converts photonic excitation into a charge separation. This subunit is found at the monomer-monomer interface. MEVNILAFIATALFILVPTAFL LIIYVKTVSQNN
Table 7: Examples of the instruction-tuning data on molecule-protein alignment.
Dataset Instruction Input Output
BindingDB Generate a drug molecule binding to the target protein. MLRQIIGQAKKHPSLIPLFVFIGT GATGATLYLLRLALFNPDVCWDRN NPEPWNKLGPNDQYKFYSVNVDYS KLKKERPDF [C][O][C][=C][C][Branch2] [Ring1][O][C][O][C][=Branch1] [C][=O][C][=C][C][=C][O][C] [Branch1][C$][C$][=C][C][Ring1][Branch2][=C] [Ring1][N][=C][C][Branch1] [Ring1][O][C][=C][Ring2] [Ring1][Branch2][O][C]
Rhea Generate an enzyme that can catalyze for the given substrate. [O][=C][Branch1][C][O-1] [C][=Branch1][C][=O][C][O] DLFHAQRGHGNLTQTLTDYMPYIGHIQI SQVPSRHEPDSDGEINYPFIFHTIAKLG YKGWVGCEYTPRGKTQLTV
Table 8: Statistics of datasets for molecule-natural language alignment and protein-natural language alignment.
Dataset Training Set Test Set
ChEBI 26,407 3,300
SwissProt_Name 455,583 3,000
SwissProt_Family 370,642 3,000
SwissProt_Location 275,740 3,000
SwissProt_Function 411,064 3,000
SwissProt_Description 393,818 3,000
Table 9: Statistics of datasets for molecule-protein alignment. # of Entries and # of Targets denote the number of data entries and the number of targets, respectively.
Dataset Training Set Test Set
# of Entries # of Targets # of Entries # of Targets
BindingDB 335,450 2,033 1,612 100
Rhea 190,206 926 2,207 100
Figure 1: Case analysis for molecule captioning task.
Figure 2: Case analysis for description-based molecule generation task.
Table 10: Case analysis for protein function answering task.
UniProtID Ground truth Generation BLEU-2 ROURGE-2
Q9NRY2 Component of the SOSS complex, a multiprotein complex that functions downstream of the MRN complex to promote DNA repair and G2/M checkpoint. The SOSS complex associates with single-stranded DNA at DNA lesions and influences diverse endpoints in the cellular DNA damage response including cell-cycle checkpoint activation, recombinational repair and maintenance of genomic stability. Required for efficient homologous recombination-dependent repair of double-strand breaks (DSBs) and ATM-dependent signaling pathways. Component of the SOSS complex, a multiprotein complex that functions downstream of the MRN complex to promote DNA repair and G2/M checkpoint. The SOSS complex associates with single-stranded DNA at DNA lesions and influences diverse endpoints in the cellular DNA damage response including cell-cycle checkpoint activation, recombinational repair and maintenance of genomic stability. Required for efficient homologous recombination-dependent repair of double-strand breaks (DSBs). 92.8 96.4
P73070 One of the components of the core complex of photosystem II (PSII), possibly involved in regulating electron flow. PSII is a light-driven water:plastoquinone oxidoreductase that uses light energy to abstract electrons from H(2)O, generating O(2) and a proton gradient subsequently used for ATP formation. It consists of a core antenna complex that captures photons, and an electron transfer chain that converts photonic excitation into a charge separation. One of the components of the core complex of photosystem II (PSII). PSII is a light-driven water:plastoquinone oxidoreductase that uses light energy to abstract electrons from H(2)O, generating O(2) and a proton gradient subsequently used for ATP formation. It consists of a core antenna complex that captures photons, and an electron transfer chain that converts photonic excitation into a charge separation. 91.7 94.1
Q9Y2G3 Catalytic component of a P4-ATPase flippase complex which catalyzes the hydrolysis of ATP coupled to the transport of aminophospholipids, phosphatidylserines (PS) and phosphatidylethanolamines (PE), from the outer to the inner leaflet of intracellular membranes. May contribute to the maintenance of membrane lipid asymmetry in endosome compartment. Catalytic component of a P4-ATPase flippase complex which catalyzes the hydrolysis of ATP coupled to the transport of aminophospholipids from the outer to the inner leaflet of various membranes and ensures the maintenance of asymmetric distribution of phospholipids. Phospholipid translocation seems also to be implicated in vesicle formation and in uptake of lipid signaling molecules. May also participate in the establishment of the thrombopoietin gradient across the membrane of platelets. 45.9 47.4
Q9FY89 Component of the ESCRT-III complex, which is required for multivesicular bodies (MVBs) formation and sorting of endosomal cargo proteins into MVBs. The ESCRT-III complex is probably involved in the concentration of MVB cargo. Probable core component of the endosomal sorting required for transport complex III (ESCRT-III) which is involved in multivesicular bodies (MVBs) formation and sorting of endosomal cargo proteins into MVBs. MVBs contain intraluminal vesicles (ILVs) that are generated by invagination and scission from the limiting membrane of the endosome and mostly are delivered to lysosomes enabling degradation of membrane proteins, such as stimulated growth factor receptors, lysosomal enzymes and lipids. 34.8 33.0
P0CP67 Responds to activation by environmental stress by phosphorylating downstream targets. Responds to activation by environmental stress and pro-inflammatory cytokines by phosphorylating a number of transcription factors, primarily components of AP-1 such as c-Jun and ATF2 and thus regulates AP-1 transcriptional activity. May play a role in the regulation of the circadian clock. 15.3 22.2
Figure 3: Case analysis for description-based protein generation task.
Figure 4: a, Performance comparison on protein-based drug discovery task. Comparing InstructBioMol and baselines in terms of the Top-1 Vina Score against all proteins in the test set. b, c, Performance comparison on substrate-based enzyme design task. Comparing InstructBioMol and baselines in terms of the Top-1 ESP Score (b) and Top-1 Vina Score (c) against all substrates in the test set.